4 Apr 0515
Sensor activations at 0500 hrs, string 156 target C, at grid BS65393544. 8 persons moving S-N, metal confirmed, Arty engaged w/10 rounds 4.2 HE,6 rounds 105 HE.
4 Apr 1145
Co B requests urgent dust off for line # 88 & # 46 men injured by booby trap at grid BS683390, dust off requested at 1144 hrs. Ref # 23, two persons are injured by booby trap:
Crabtree, David D. Sp/4 -fragment wounds in the left leg. Busby, Stephen, L. Pfc -multiple fragment wounds to the front of the body, 1 hole in front of throat, dust off complete at 1157 hrs.
4 Apr 1300
Ref # 23 Busby died of wounds at the 23d Med in Duc Pho.
4 Apr 1540
Sensor activation string # 156, 8 persons stationary confirmed by metal, 4.2 expended 10 rounds HE...Results unknown.
5 Apr 1816
Dust off requested by Co C .....Jones Jerry, Sp/4 gash in hand and dust off was complete at 1832 hrs.
5 Apr 2005
Spot Report: At 1630 hrs at grid BS713413, Co B found one 33 H/G rigged as booby trap located on the side of a trail, the grenade was on the ground w/ a pressure release device, it was blown in place, diagram to be submitted.
6 Apr 1415
At 1345 hrs Co A 3rd platoon at grid BS719349. heard movement ahead of them, they then proceeded and found 1 satchel charge (size of a # 10 can) hanging from a tree being set up as a booby trap (not completed) and also found 2 spider holes being dug along trail.
7 Apr 1100
Spot Report: At 1053 hrs. Co B found at grid BS702394, a mechanical mouse trap ambushed in a old NDP, the trap was sprung with a trip wire removed, claymore reversed, evidence of explosion under claymore, it was blown in place.
7 Apr 1104
Co D found grid at BS676413. 1 dud 250 pound bomb, requesting a EOD Team to dispose.
7 Apr 1115
Spot Report: At 1015 hrs. Co A 3rd platoon hit a booby trap at grid BS717343, possibly a hand granade set up with a pressure device, results: 2 US WIA, dust off was requested at 1018 hrs. Bde & Bn S-1 notified. At 1020 hrs, dust off E-5 Beaver, Donald A. & Mills, Donald, J. E-3 complete at 1055 hrs.
8 Apr 1027
081010 Apr coord at grid BS714342 2nd platoon Co A found 3 ...6' x 10' hooches, 2....8' x 10' hooches, 2 tunnels (bomb shelter type) 1 chicom, 30 pistol rounds, 1 pistol magazine (cal unk) 3 rounds 9mm, 3 m-16 magazines, 62 M-16 rounds, 200 pungi states, age of hooches is 3-4 months old...last used recently-1 day ago. All military structures were destroyed.
8 Apr 1235
Sensor activation at 1215 hrs, string 152 A at grid BS65853761, 3 persons moving S confirmed by heat, engaged w/10 rounds 4.2 HE.....Results: unknown.
9 Apr 0820
Message from 1-20 Inf: On 8 Apr engaged 10 VC/NVA at grid BS 724440, results: 1 CPT KIA, documents revealed that the CPT had 20 persons at base of a mountain location and name unknown, documents also revealed the unit was to hit LZ Liz in the next 10 days. 1-20 thinks the base camp may be located in and around the 515 valley in our AO.
9 Apr 1130
Received from Co B at 1100 hrs at grid BS718375: Co B engaged 1 VC who was the point for several persons, results: 1 NVA WIA, captured 1 AK-44 and magazine, 4 rucksacks, 3 H/G's, 1 baseball frag, 2 old grenades and 3 maps.
9 Apr 1145
Ref # 36, additional information: Co B engaged 4 VC, 1 WIA was point, 2nd person opened w/1 magazine AK and evaded, Co B is pursuing at this time.
9 Apr 1525
Ref # 36, Co B found a blood trail approx 50 meters from earlier contact, grid BS718375 which they followed until it stopped. Co is moving down same trail.
9 Apr 1220
Sensor activation at 1210 hrs, a string 153 B at grid BS657410 sensed 4 persons moving N-S, confirmed by heat, engaged w/10 rounds 4.2 HE, results: unknown.
9 Apr 1530
Spot Report: At 1507 hrs Dolphin 19 observed 3 persons at grid BS670403 moving near old NDP, engaged w/M-60 from doorgunner and 4.2-15 rounds from LZSJ, Results unknown.
10 Apr 0945
Sensor activation on 4-21st Inf string # 74 at grid BS875329 4 persons moving N-S confirmed by metal.
10 Apr 1401
Spot Report: from Co B at grid BS701387 Co B found 1 hut 10' x 10' used for storage...hut used last 1 month ago. des -troyed.
11 Apr 1350
Co B 1st platoon at grid BS671413. Found 3 trails 1 from East to West, 1 from NE to above grid. Observed signs of 3-5 persons moving West probably this morning.
11 Apr 1620
From Bde, LRRP Idaho at grid BS692485 had 1 VC in black pj's...came up from rear and engaged w/AK-47, LRRP's returned fire and enemy evaded to the West. LRRP Team heard 3 warning shots from NW, W, and SW. Negative casualties.
12 Apr 1215
At 1215 hrs. Apr at grid BS701355 Co A 2nd platoon engaged with 7 VC. They have 1 VC KIA, also 1 AK-47 CIA, more to follow.
12 Apr 1305
Ref # 31, Co A 2nd platoon element was walking down a trail when the point man w/element engaged 4 VC at 35 meters and the rear element engaged 2 VC at 30 meters, 1 unit VC was heading South, 1 unit was going NE, results...1 VC KIA, 1 AK-47 CIA #10080446 w/10 round clip full in weapon, on KIA was 1/2 lb of C-4, 1 tube of tootpaste and 1 pair of GI pants, terrain was jungle.
12 Apr 2011
Co B Spot Report: 3rd platoon was setting up a night ambush security element when it spotted 4 VC/NVA moving West to East along 515 rounds via BS692419. EM spotted patrol and started evading. Security engaged w/ S/A, unable to engage w/Arty due GT line. It was too late to see. EM appeared to be armed and wearing uniform. Will check out area at first light.
12 Apr 2235
Radar at 2225 hrs. sighted 4 to 5 persons moving W azmith 5440 dist 257 meters also another group at azmith 5280 dist 2620 covering at az 5440 dist 2620 8-12 persons w/ 4.2 ...Results unknown.
13 Apr 0345
At 0330 hrs. 3 persons were spotted at bunker #'s 2-3-4 area, 1 with weapon, possibly and another with a larger instument, engaged w/ Arty and 81's.... 9 rounds 105.
At 1315 hrs. Sensor activation String 153 at grid BS660411...Sensed 3 persons moving North confirmed by heat. Target was initially moving at 3 mph, and reversed direction and activated sensors 153 B, C after 153A was initially, fired 6 105 HE rounds , & 30 rounds of 4.2..with results unknown.
At 1335 hrs. sensor activation string B at grid BS 657410 sensed 5 persons moving S confirmed by heat, rate of travel not given, engaged w/10 rounds 4.2 6 rounds 105 HE....results unknown.
Radar 0058 at grid BS65393544 string 156 target C 4...persons moving S-N metal confirmed, engaged w/4 rounds 105 & 10 rounds 4.2 HE...results unknown.
Grid BS702355 at 1135 hrs. 1st platoon Co A heard voice in the brush in the vicinity of the above grid, engaged w/s/a fire M-79's, and 81mm, moving into area at this time.
Ref # 42, follow up information, they proceeded to check out area, spotted persons evading and egaged persons, they found the persons were a 6 year old VN hit in hip, and dust off called in urgent....complete 1245 hrs.
Recon T1 142010 spotted a camp fire, engaged w/4.2....12 rounds HE-had one large secondary explosion.
Co C while OPCON to 4-21st Infantry; hit unknown type booby trap, results; 6 US KIA...13 US WIA;
TMF at 2030 hrs. at grid BS61793888 string 154 target 154 B metal confirmation 15-20 persons, engaged w/12 rounds HE. Had movement and engaged 10 more rounds HE.
Received Weather Report frm Sgt Campbell...Weather thru 16 Apr temp 81-69, humid-95-65, wind direction NE 15 knots and outlook mostly cloudly 80 % illum.
Alpha 4/3 11th